LOWEST ENERGY VIBRATIONAL MODES OF NINE NAPHTHALENE DERIVATIVES; EXPERIMENT AND THEORY

\underline{M. A. MARTIN-DRUMEL}^{a}, O. PIRALI, Y. LAQUAIS^{b}, Institut des Sciences Moléculaires d’Orsay, CNRS, UMR 8214, Université Paris XI, bât. 210, 91405 Orsay Cedex, France; SOLEIL Synchrotron, AILES beamline, L’orme des Merisiers, Saint-Aubin, 91192 Gif-Sur-Yvette, France; C. FALVO, P. PARNEIX and PH. BRECHIGNAC, Institut des Sciences Moléculaires d’Orsay, CNRS, UMR 8214, Université Paris XI, bât. 210, 91405 Orsay Cedex, France.

FIR gas phase absorption spectra of nine naphthalene derivatives (azulene, quinoline, isoquinoline, biphenyl, diphenylmethane, bibenzyl, 2-, 3-, and 4-phenyltoluene) have been recorded at medium resolution (0.5 cm^{−1}) using a Fourier transform Bruker IFS125 interferometer located on the AILES beamline of SOLEIL synchrotron. Assignments of these weak vibrational bands were performed using density-functional theory calculations carried out at the harmonic and anharmonic levels (B97-1/6-311G(d,p)). Molecular structure dependence of the FIR spectra is discussed based on the comparative study of several selected FIR modes.

aPresent address: LPCA, EA 4493, Université du Littoral Côte d’Opale, 59140 Dunkerque, France

bPresent address: CEA, IRAMIS, SPAM, Lab. Francis Perrin, 91192 Gif-sur-Yvette, France